OpenStudy (anonymous):

solve each system by substitution. 2x+5y=22 y=-3x-6

OpenStudy (anonymous):

2x + 5(-3x - 6) = 22 2x - 15x - 30 = 22 -13x = 52 x = -4 y = -3(-4) - 6 y = 12 - 6 y = 6 Not sure if you needed y.